# Instructions for *ACL Proceedings Learning to Tackle and Generalize to TurkingBench Can Self-Supervised Models Learn to Tackle TurkingBench?

### v3

@@ Abstract
Building on this idea, we present TurkingBench, a benchmark consisting of tasks presented as web pages with textual instructions and multi-modal contexts. Unlike previous approaches that rely on artificially synthesized web pages, our benchmark uses natural HTML pages originally designed for crowdsourcing workers to perform various annotation tasks. Each task’s HTML instructions are instantiated with different values derived from crowdsourcing tasks, creating diverse instances. This benchmark includes 32.2K instances spread across 158 tasks.
To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, selecting a radio button). We assess the performance of cutting-edge private and open-source models, including language-only and vision-language models (such as GPT4 and InternVL), on this benchmark. Our results show that while these models outperform random chance, there is still significant room for improvement. We hope that this benchmark will drive progress in the evaluation and development of web-based agents.111Dataset and code: https://turkingbench.github.io
@@ 1 Introduction
We present TurkingBench, a benchmark for web-grounded tasks. Examples of these tasks can be seen in Figure 1. The information in these tasks is conveyed through a mix of stylistic elements (such as colors, font sizes, and shapes), structural features (like tables and paragraph headings), and multiple modalities (including text, images, and videos). Effectively interpreting these multi-modal signals is a complex challenge that demands a range of skills, particularly the ability to understand language within rich visual contexts.
The tasks in TurkingBench are drawn from those published on the Amazon Mechanical Turk (AMT) platform, a well-known site where human workers complete micro-tasks for rewards. The benchmark comprises 158 web-grounded tasks (web pages), each averaging 16.8K tokens and containing 15.6 input fields that AI models must label or fill in. Each task can be instantiated with different input values (e.g., varying questions, paragraphs, images). In total, our benchmark includes 36.2K input fields. There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see”, “read” or “head”. (ii) Interaction—agents’ actions must consider the context of previous decisions on the web pages. (iii) Long context—web pages (HTML, screenshots, etc.) often are quite long.
@@
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. Among the open-source models, InternVL-2 achieves the highest performance. In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches. Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset. We analyze model performance based on field type and task length, identifying challenges for future progress.
In summary, we present TurkingBench, a benchmark for web-based tasks requiring multi-modal and interactive reasoning. We establish an evaluation framework for interacting with and modifying web pages. Our evaluation of notable models reveals significant room for improvement. We hope this benchmark motivates and aids the advancement of assistive, web-based agents.
@@ 2 Related Work
Unlike our focus on web pages, it is worth noting the benchmarks that utilize other environments for interactive problem solving. For instance, benchmarks that concentrate on mobile operating systems such as Android Li et al. (2020); Toyama et al. (2021); Sun et al. (2022); Burns et al. (2022) or computer applications in0 operating system environments Trivedi et al. (2024); Xie et al. (2024).
Multi-modal reasoning tasks.
Web-based agents.
@@ 3.1 Benchmark Schema
An example is shown is shown in Figure 2. As the example shows, the HTML template contains variables (e.g., ${sys10}) that are instantiated with input values. Furthermore, the HTML template contains input fields for receiving input values from the web page users. For each field, we have values previously collected from crowd workers that we will use for evaluation.
@@ 3.2 Collecting the web-based tasks
3.3 Statistics
Table 2 shows the overall statistics of TurkingBench. The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Furthermore, Figure 3 shows the distribution of the input fields represented in our benchmark. The most common field is “radio,” which is expected as it is commonly used to identify preferences. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
@@ 3.4.1 A Library of Web Actions
The overall task can be viewed as iterative tool resolution based on the context of each task. A web agent solving this task must use a sequence of actions from our action library, guided by the task’s instructions (see an example in Figure 5). This approach connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), which use tools to ground LM generations in physical environments. The actions in our library serve as the “tools” that can be executed to achieve specific outcomes. The core of this setup is an LM aware of the tools, the problem context, and the ultimate goal.
@@ 3.6 Evaluation Protocol
Our evaluation protocol consistent of a loop over the various choices of evaluation tasks. Specifically means that, the evaluation script provides the model with the URL where the task can be reached. Furthermore, we provide the model the access to our action library that contains various actions that can be executed by the model.
@@ 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
Models.
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024. For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al. (2024). Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
Encoding the tasks for evaluation.
When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving. To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks. This adds up to web pages evaluated with a total of roughly input fields evaluated.
@@
We implement an oracle baseline that mimics the gold labels for each input. This oracle agent ensures the functional correctness of our evaluation. Our end-to-end evaluation process (model output execution lookup answers from web pages evaluation) is significantly more complex than a typical NLP benchmark, and any of these steps can fail. Achieving 100% accuracy with the oracle (ensuring functional correctness) took the lead student several months of effort. The oracle baseline essentially replicates the action sequences of crowdworkers for each HTML input element in the task web pages, executing appropriate actions similar to those shown in Figure 5. A pseudo-code of the oracle model is shown in Pseudocode 2.
@@ 4.1 Empirical Results
Llama3.1-Instruct (8B params) notably outperforms GPT4 (text-only), achieving a score of 25.0% compared to GPT4’s 21.3% with 7 demonstrations when “relevant” bits of the input HTML are supplied. This result is particularly impressive given that Llama3.1-Instruct operates with significantly fewer parameters (8B) than GPT4’s rumored size. The comparison highlights Llama3.1-Instruct’s ability to efficiently leverage the provided context, potentially making it better suited for certain tasks despite its smaller size. Additionally, InternVL2 demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML. InternVL2’s score of 31.0% is remarkably close to GPT4’s 33%, despite the substantial difference in parameter sizes. This near-equivalence underscores the potential of InternVL2 to compete with proprietary models, offering similar performance with a more compact architecture.
Other open-source models, such as LLaVA-VL-Vicuna/Mistral (7B and 13B parameters), show varying degrees of success, with performance scores ranging from 13.8% to 22.7%. While these models are still behind Llama3.1 and InternVL2, their results demonstrate the diverse capabilities and potential of open-source approaches.
Text-only and vision-language models show complementary capability.
@@
For our top configurations in section 4 (GPT4, 7 demonstrations with “full” HTML encoding) we show the breakdown of the model performances across various tasks in Figure 6. On 9 evaluation tasks (out of 16) the performance of these two models are notably different from one another. While it is intuitive to speculate that task with richer task interface likely benefit more from visual, empirical this it was not obvious what exact details details about task design leads to such differences.
The gains of in-context demonstrations plateau quickly.
@@
As mentioned, we use few-shot prompting of models in order to steer them their predictions. Here we study the effect of the number of demonstrations in model performance. As the results are shown in Figure 7, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
Performance across field types.
@@
To better understand the task difficulty based on field types, we show the performance of GPT4 (text input; 7 demonstrations) for each input field type in Figure 8. We distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs. <textarea>). The former is typically used for short inputs, while the latter is for longer, multi-sentence outputs. A notable finding was the surprisingly lower performance of GPT4 compared to InternVL2 on “text” input fields, which we could not explain. However, since the data is skewed toward checkboxes (Figure 3), GPT4 performs better in the aggregate evaluations.
@@ 4.2 Error Analysis
Another significant discrepancy was observed in tasks requiring GPT4o to generate both correct and incorrect answers. Instead of providing actual incorrect answers, GPT4o sometimes returned placeholders like //incorrect options//, failing to meet the task’s requirements. In some cases, GPT4o also made syntax errors. However, the evaluation focused on the content of the responses rather than their syntactical correctness, so syntax errors did not affect the assessment of answer accuracy. We would like to provide more details upon having more space.
@@ 5 Conclusion and Future Work
Limitations.
We discuss several limitations: (i) No navigation between pages: Our benchmark does not require navigation between web pages. While inter-page navigation is important, it is not the only challenge. Effective understanding and manipulation of each page, which is our focus, remains a significant challenge for web agents. Our benchmark still requires navigation within each page, involving multiple rounds of interaction to solve tasks with multiple input fields. As shown in our experiments, this remains challenging for the models. We accept this trade-off to obtain more natural tasks compared to most existing benchmarks. (ii) Simplified evaluation: This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation. For simplicity, our evaluation setup provides some guidance by specifying the names of the fields to be modified. Future work should explore variants of our experiments where such hints are not provided to the model.
Ethical considerations.

### v4

# Tur[k]ingBench: A Challenge Benchmark for Web Agents

@@ Abstract
Building on this idea, we present TurkingBench, a benchmark consisting of tasks presented as web pages with textual instructions and multi-modal contexts. Unlike previous approaches that rely on artificially synthesized web pages, our benchmark uses natural HTML pages originally designed for crowdsourcing workers to perform various annotation tasks. Each task’s HTML instructions are instantiated with different values derived from crowdsourcing tasks, creating diverse instances. This benchmark includes 32.2K instances spread across 158 tasks. To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, or selecting a radio button). We assess the performance of cutting-edge private and open-source models, including language-only and vision-language models (such as GPT4 and InternVL), on this benchmark. Our results show that while these models outperform random chance, there is still significant room for improvement. We hope that this benchmark will drive progress in the evaluation and development of web agents.111 Data and code: https://turkingbench.github.io/
NAACL ’25
Tur[k]ingBench: A Challenge Benchmark for Web Agents
Kevin Xu∗ Yeganeh Kordi Tanay Nayak Adi Asija Yizhong Wang Kate Sanders Adam Byerly Jingyu Zhang Benjamin Van Durme Daniel Khashabi††thanks: Correspondence to: {kxu39,danielk}@jhu.edu Johns Hopkins University Brown University University of Washington
@@ 1 Introduction
We present TurkingBench, a benchmark for web-grounded tasks. An example task can be seen in Figure 1. The information in these tasks is conveyed through a mix of stylistic elements (such as colors, font sizes, and shapes), structural features (like tables and paragraph headings), and multiple modalities (including text, images, and videos). Effectively interpreting these multi-modal signals is a complex challenge that demands a range of skills, particularly the ability to understand language within rich visual contexts.
The tasks in TurkingBench are drawn from those published on the Amazon Mechanical Turk (AMT) platform, a well-known site where human workers complete micro-tasks for rewards. The benchmark comprises 158 web-grounded tasks (web pages), each averaging 16.8K tokens and containing 15.6 input fields that AI models must label or fill in. Each task can be instantiated with different input values (e.g., varying questions, paragraphs, images). In total, our benchmark includes 36.2K input fields. There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see” or “read”. (ii) Interaction—agents’ actions must consider the context of previous decisions on the web pages. (iii) Long context—web pages (HTML, screenshots, etc.) often are quite long.
@@
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. Among the open-source models, Qwen2 achieves the highest performance. In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches. Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset. We analyze model performance based on field type and task length, identifying challenges for future progress.
We chose crowdsourcing tasks for their diverse web interactions, including multimodal reasoning, complex form filling, and sequential dependencies—common in e-commerce, education, and administration. While rooted in this domain, our benchmark’s challenges (e.g., diverse inputs, long-context reasoning, and multimodal interactions) apply broadly. Effective page understanding and manipulation remain key challenges for web agents.
@@
In summary, we introduce TurkingBench, a benchmark for multimodal, interactive web-based tasks, along with an evaluation framework for web interaction. Our analysis of leading models highlights significant gaps, aiming to drive progress in assistive web agents.
@@ 2 Related Work
Unlike our focus on web pages, it is worth noting the benchmarks that utilize other environments for interactive problem solving. For instance, benchmarks that concentrate on mobile operating systems such as Android Li et al. (2020); Toyama et al. (2021); Sun et al. (2022); Burns et al. (2022) or computer applications in operating system environments Trivedi et al. (2024); Xie et al. (2024).
@@ 3.1 Benchmark Schema
An example is shown in Figure 2. As the example shows, the HTML template contains variables (e.g., ${sys10}) that are instantiated with input values. Furthermore, the HTML template contains input fields for receiving input values from the web page users. For each field, we have values previously collected from crowd workers that we will use for evaluation.
@@ 3.2 Collecting the web-based tasks
Statistics.
The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Table 2 shows the overall statistics of TurkingBench. Furthermore, Figure 3 shows the distribution of the input fields represented in our benchmark. The most common field is “checkbox,” which is expected as it is commonly used when multiple options could be true. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
@@ 3.3.1 A Library of Web Actions
The task here can be framed as iterative tool resolution based on task context. A web agent must execute a sequence of executable actions from our library, guided by task instructions (see Figure 1). This aligns with tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), where tools ground LM generations in an environment. Here, our library’s actions function as “tools” for achieving specific outcomes, with an LM aware of the tools, task context, and final goal.
@@ 3.5 Evaluation Protocol
Our evaluation iterates through different choices of evaluation tasks (Pseudocode 1). Specifically, the evaluation script supplies the model with the URL where the task is accessible. Additionally, the model has access to our action library, enabling executable actions.
@@
Screen Size Considerations.
When benchmarking multimodal models, the capture_screen action retrieves visual information from the screen. A potential concern is the impact of screen size on benchmarking consistency. To address this, all evaluations using capture_screen run in headless mode, where the browser operates without a monitor and uses predefined, consistent virtual screen dimensions. This eliminates hardware-specific influences, ensuring robust and reproducible results unaffected by screen size variability.
@@ 4 Evaluating Models in Solving Web-based Tasks in TurkingBench
Models.
As discussed earlier, our models need to consume information on a mix of information modalities. A model can, therefore, consume the instructions as text, image, or a combination of both. We experiment with state-of-the-art proprietary models like GPT4 and Claude-2.1.666GPT4/GPT4-V/Claude accessed via APIs in January 2024; GPT4o accessed in August 2024. For GPT4 models, we experiment with both the text models and the vision-language variant that is trained in a joint of visio-textual information. We compare the performance of these models with the latest open-source vision language models like LLaVA-1.6 (Liu et al., 2023a), InternVL2 (Chen et al., 2024) and text-only models like Llama-3.1-Instruct Dubey et al. (2024) and Qwen2 Yang et al. (2024). Our evaluation focused on major model families (GPT4, LLaVA1.6, etc.), as our goal was setting baselines with widely-recognized models.
Encoding the tasks for evaluation.
When processing the HTML content of the tasks, we consider two variants: (1) “full” indicates all the HTML content of the entire page. (2) “relevant” indicates a few neighboring lines of HTML adjacent to (above and under) the input field the model is currently solving (further details in §C.) To reduce the cost of evaluation, we use 20 instances per each of the 20 evaluation tasks. This adds up to web pages evaluated with a total of roughly input fields evaluated.
@@
We implement an oracle baseline that mimics the gold labels for each input (see the logic in 2). 777See the a screencast of the oracle’s execution. This oracle agent ensures the functional correctness of our evaluation. Our end-to-end evaluation process (model output execution lookup answers from web pages evaluation) is significantly more complex than a typical NLP benchmark, and any of these steps can fail. Achieving 100% accuracy with the oracle (ensuring functional correctness) took the lead student several months of effort. The oracle baseline essentially replicates the action sequences of crowdworkers for each HTML input element in the task web pages, executing appropriate actions similar to those shown in Figure 1.
@@ 4.1 Empirical Results
Llama3.1-Instruct (8B params) notably outperforms GPT4 (text-only), achieving a score of 25.0% compared to GPT4’s 21.3% with 7 demonstrations when “relevant” bits of the input HTML are supplied. This result is particularly impressive given that Llama3.1-Instruct operates with significantly fewer parameters (8B) than GPT4’s rumored size. The comparison highlights Llama3.1-Instruct’s ability to efficiently leverage the provided context, potentially making it better suited for certain tasks despite its smaller size. Additionally, Qwen2 (72B) demonstrates strong performance, coming very close to GPT4 when evaluated with the “relevant” HTML. Qwen2’s score of 34.1% is not too far off from GPT4-V’s 41.7%, which underscores the potential and promise of open-source models over proprietary models.
4.2 Analysis: Test vs Vision Models
@@
For one of our top configurations in section 4 (GPT-4, 7 demonstrations with “full” HTML encoding), we present a breakdown of model performance across tasks in Figure 5. On 9 out of 16 evaluation tasks, the two models exhibit notable performance differences. While it’s intuitive to assume tasks with richer interfaces benefit more from visual input, empirical results do not clearly pinpoint which task design aspects drive these differences. Overall, we conclude that text-only and vision-language models have complementary capabilities.
4.3 Analysis: Effect of the Number of Demos
@@
As mentioned earlier, we use few-shot prompting of models in order to steer them their predictions. Here we study the effect of the number of demonstrations in model performance. As the results are shown in Figure 6, the gains of in-context demonstrations quickly plateaus when the number of demonstrations just above 3 demonstrations.
4.4 Analysis: Performance Per Field Types
@@
We provide extended details on performances per input field type in Table 5. Note that we distinguish between "text" and "textarea" fields because they are defined with different HTML tags (<input type="text"> vs. <textarea>). The former is typically used for short inputs, while the latter is for longer, multi-sentence outputs. There is no consistent patter here. Most models tend to have similar weaknesses and strengths, which is not surprising as they’re generally trained on similar-ish [pre-training] data. A notable finding was the surprisingly lower performance of GPT4 compared to other models on “text” input fields, which we could not explain. However, since the data is skewed toward checkboxes (Figure 3), GPT4 performs better in the aggregate evaluations. This suggests the potential for hybrid models by combining those that excel at specific input field types.
Model | # of params | # of demos | Modality | all | checkbox | radio | text | textarea |
Llama-3.2 | 3 B | 3 | T | 20.8 | 58.7 | 25.8 | 19.1 | 0.3 |
Llama-3.2 | 3 B | 7 | T | 20.3 | 56.8 | 22.9 | 22.0 | 2.9 |
Gemma-2 | 7 B | 3 | T | 20.4 | 54.1 | 28.7 | 15.1 | 2.0 |
InternVL | 40 B | 3 | T+ V | 30.7 | 59.0 | 85.0 | 46.0 | 12.0 |
InternVL | 40 B | 7 | T+ V | 26.1 | 85.0 | 28.2 | 48.1 | 7.5 |
InternVL | 76 B | 3 | T+ V | 30.4 | 63.7 | 29.7 | 48.4 | 17.1 |
InternVL | 76 B | 7 | T+ V | 29.6 | 56.6 | 30.5 | 46.4 | 12.3 |
GPT4-VL | - | 3 | T+ V | 30.3 | 73.3 | 34.2 | 16.4 | 15.5 |
@@ 4.5 Error Analysis
Another significant discrepancy was observed in tasks requiring GPT4o to generate both correct and incorrect answers. Instead of providing actual incorrect answers, GPT4o sometimes returned placeholders like //incorrect options//, failing to meet the task’s requirements. In some cases, GPT4o also made syntax errors. However, the evaluation focused on the content of the responses rather than their syntactical correctness, so syntax errors did not affect the assessment of answer accuracy. In Table 6 we highlight examples of model predictions for a given UI.
@@ 5 Conclusion and Future Work
Limitations
We discuss several limitations:
Intra page navigation.
Our benchmark does not require navigation between web pages. While inter-page navigation is important, it is not the only challenge. Effective understanding and manipulation of each page, which is our focus, remains a significant challenge for web agents. Our benchmark still requires navigation within each page. This is because almost all of our tasks have more than one step (input field), each needed to be solved in a different round of interaction. As shown in our experiments, the benchmark remains challenging for the models. We accept this trade-off to obtain more natural tasks compared to most existing benchmarks.
Simplified evaluation.
This is not an inherent limitation of our benchmark but a simplifying assumption in our evaluation. For simplicity, our evaluation setup provides some guidance by specifying the names of the fields to be modified. Future work should explore variants of our experiments where such hints are not provided to the model.
Scope.
Our benchmark’s distribution is biased to crowd-sourcing tasks and hence it is not a holistic measure of web-based agents. That being said, all benchmarks carry their own biases. In the context of web-based tasks, all the existing benchmarks (WebShop, Mind2Web, WebArena, etc.) have their own set of assumptions and restrictions about the scope of web-based agents. We view all these efforts to be complementary, each quantifying a unique aspect of intelligent behavior on the web.
Ethical Considerations
@@ Acknowledgements
Supplemental Material
Appendix A Additional Related Work
Here we discuss other related work that did not fit in the main text.
Multi-modal reasoning tasks.
Web-based agents.
Appendix B Pseudocode for the oracle baseline
Appendix C Extracting the “relevant” HTML code for each field
We detail the procedure for retrieving "relevant" neighboring HTML, as discussed in §4 under "Encoding the tasks for evaluation." The “relevant” HTML encoding is implemented by extracting and returning a subset of HTML adjacent to (above and under) the specific input field on a webpage.Specifically: (1) We retrieve the full HTML content of the target elements’ grandparents elements. This potentially contains a lot of content. (2.) We split this code into html elements (split based on occurrence of ">") (3) We select 15 elements before and 30 elements below the target element. The original function is accessible in our public code base.
Category: Both automatic and human evaluations classified the action as incorrect (true negative). |
Explanation: It is perfectly legal to do carry your pet in public, even as a status symbol. |
Category: Both automatic and human evaluations classified the action as incorrect (true negative). |
Explanation: Our society does not "strongly pressure" people to carry pets in public. |
Category: Automatic evaluation rated the action as incorrect but human evaluations rated it correct (false negative). |
Explanation: The automatic evaluation has rated it incorrect since ‘previous’ was not found in the list of gold answers: [’filled’, ’this’, ’last’]. |
Category: Automatic evaluation rated the action as incorrect but human evaluations rated it correct (false negative). |
Explanation: The automatic evaluation has rated it incorrect since ‘great’ was not found in the list of gold answers: [’good’, ’guy’, ’girl’, ’Good’, ’big’]. |
Category: Automatic evaluation and human evaluation both rated the action as correct (true positive). |
Explanation: – |
Category: Automatic evaluation and human evaluation both rated the action as correct (true positive). |
Explanation: – |
Category: Automatic evaluation rated the action as correct even though human evaluations rated it incorrect (false positive). |
Explanation: The statement is true, but the highlighted region is irrelevant. |
Category: The model response does not adhere to the expected syntax. |
Explanation: The execution here failed since instead of selecting one of the radio values from -2 to 2, the model responded with a blank string. |